# Using Machine Learning to Explore Defect Configurations in Cd/Zn-Se/Te Compounds

### v1

@@ 2 Description of the DFT datasets
For the GGA-PBE dataset, we collected a majority of the structures and energies from multiple past published works from our group 82, 83, 29, 45, and combined them with some systematic new computations on previously unexplored alloy compositions and defect structures. Upwards of 20,000 bulk and defect configurations were part of this dataset, along with their total energies, atomic forces, and stresses collected from the DFT geometry optimization trajectories. Starting from selected bulk and defect configurations from the PBE dataset, HSE06 geometry optimization was then performed using a default mixing parameter of $\alpha=0.25$. For defect calculations, lattice parameters were changed from the PBE-optimized values to the correct HSE06-optimized values to enable relaxation with fixed volume.
Combining the bulk structures (binary, ternary, and quaternary compounds; specific alloy series are listed in Figure 2(a)) and defect-containing structures into one dataset for training the GNN models helps simultaneously take into account the effects of alloying (e.g., Cd-Zn mixing or Se-Te mixing), cation and anion ordering, supercell size, and the type and distribution of point defects. Every structure is characterized by its “crystal formation energy” (CFE), which is simply the per-atom energy required to split any bulk or defect crystal into its constituent atoms. GNN models can be trained to either directly predict the CFE or indirectly predict it following a force field-based geometry optimization. From the CFE, either the bulk stability (in terms of a formation energy, decomposition energy, or energy above hull) or the defect formation energy can subsequently be estimated by using known energies of relevant reference phases.
Our chemical space includes all possible Cd/Zn-Te/Se/S binary compounds as well as selected series of ternary and quaternary compounds, specifically CdSexTe1-x, CdSxSe1-x, CdxZn1-xS, CdxZn1-xSe, CdxZn1-xTe, ZnSexTe1-x, ZnSxSe1-x, Cd0.5Zn0.5SxSe1-x, and Cd0.5Zn0.5SexTe1-x. The mixing fraction x systematically varies in multiples of 0.125 (or n/8 mixing, where n is a positive integer between 1 and 8), leading to a total of 81 unique compositions 82. All such structures are simulated in both $\times2$$\times$2 (64 atoms) and $\times3$$\times$3 (216 atoms) cubic ZB supercells using the special quasirandom structures (SQS) 84 approach, starting from known binary compound ground state configurations. From across the set of all PBE calculations, a total of 10,080 bulk configurations were collected, including intermediate structures from the geometry optimization trajectories. The CFE is calculated as follows, taking CdSexTe1-x as an example:
$CFE=\frac{E(\text{Cd${}_{a}$Se${}_{b}$Te${}_{c}$})-aE(\text{Cd})-bE(\text{Se})-cE(\text{Te})}{N_{\text{atoms}}}$ (1)
The dataset compiled from past publications contains a whole host of native defects (vacancies, self-interstitials, anti-site substitutions), extrinsic interstitials and substitutional defects accounting for both unintentional impurities and intentional doping, and a handful of defect complexes, in different Cd/Zn-Te/Se/S compositions. The number of $\times2$$\times$2 supercell defect structures for different charge states (total charge q imposed on the defect simulation during geometry optimization) accounted for the following: 7302 (q=+2), 6201 (q=+1), 8203 (q=0), 6361 (q=-1), and 7689 (q=-2). We performed a series of new defect (native and extrinsic) calculations using $\times3$$\times$3 supercell structures for some selected compounds such as CdTe and CdSe0.25Te0.75, as well as in CdTe-ZnTe heterostructures and CdTe grain boundary structures collected from other past work 85. Defect species included both native defects and extrinsic impurities and dopants of interest such as Cl, O, Cu, As, and P.
The CdTe-ZnTe interface was simulated by merging optimized $3\times 3\times 3$ supercells of CdTe and ZnTe with slightly strained lattice parameters, followed by volume-free geometry optimization. For all new defect calculations in $\times3$$\times$3 supercell binary and alloyed structures, heterostructures, and grain boundaries, the Doped package 86 was used to generate a series of starting configurations based on bond distortions and atomic rattling, to adequately account for symmetry-breaking around defect centers. Ultimately, hundreds of new defect structures were generated using larger bulk supercells and the interface and grain boundary structures, and merged with the $\times2$$\times$2 supercell defect structure data compiled from past work.
In the compiled dataset, the label bulk denotes pristine supercells without any point defects, representing ideal defect-free crystal structures. The label defect refers to supercells containing a point defect, such as a vacancy, self- or extrinsic interstitial, anti-site or extrinsic substitution, or a defect complex combining two or three of these individual defects. The label interface refers to large heterostructures with a variety of point defects at the CdTe-ZnTe interface, capturing the influence of interfacial environments on defect behavior. The label GB designates grain boundary structures containing native or extrinsic defects, which helps study the interplay between point defects and extended structural features. The statistics of the compiled PBE dataset corresponding to charge states q = +2, +1, 0, -1, and -2 are presented in Figure 2(a-g), Figure S1 and Table 1. Figure 2(d) shows violin plots capturing the distribution of CFEs (q=0) across the dataset for bulk and defect configurations (including interface and GB structures), with values ranging from $\sim$ -800 meV/atom to $\sim$ -320 meV/atom. Before GNN models were trained, energetically and configurationally similar structures were removed, leading to a reduced and curated dataset of approximately 6,923 bulk configurations and 5,910 (q=+2), 4,929 (q=+1), 6,242 (q=0), 4,795 (q=-1), and 5,666 (q=-2) defect configurations. In this work, we used two established GNN approaches to train predictive models on the DFT dataset: Atomistic Line Graph Neural Network (ALIGNN) 87, and the Materials 3-body Graph Network (M3GNet)-based machine learning force field (MLFF) 44.
3 Graph neural network models for direct prediction of crystal formation energy
ALIGNN was developed by Choudhary et al.87 and considers both two-body (bond lengths) and three-body interactions (bond angles). ALIGNN leverages both graph convolution layers and line graph convolution layers to capture short-range and long-range correlations in the crystal. For training the ALIGNN models, the learning rate was set to 0.001, an AdamW optimizer was used to update the weights and biases of the model, 4 graph convolution layers and 4 line graph layers were implemented, the cutoff radius was set to 6 Å with 12 nearest neighbors to create the crystal graph, and models were trained up to 90 epochs with a batch size of 8. We experimented with different training-validation-test splits of the dataset and found that the 60:20:20 ratio works the best.
ALIGNN models were trained to predict the CFE from any given bulk or defect crystal structure, using only the neutral charge state structures at this stage. Parity plots capturing the performance of the optimized models are presented in Figure 3, in terms of ALIGNN-predicted CFE vs DFT-computed CFE for only the test set data points. Models pictured in Figure 3(a-c) are respectively trained only on bulk structures, only on defect structures, and on both bulk and defect structures; this distinction is made to understand how sensitive the models are to different types of configuration. As shown in Figure 3(a), the ALIGNN model for bulk structures alone shows a test prediction root mean squared error (RMSE) of 1.43 meV/atom, and this error remains practically unchanged for the combined model in Figure 3(c). For the defect-only ALIGNN model in Figure 3(b), test RMSE ranges from 3.09 meV/atom for interstitial defects to 4.87 meV/atom for substitutional defects to 8.36 meV/atom for vacancy defects. Each of these defect prediction errors comes down for the combined data model, proving the value of increasing the size and chemical and structural diversity of the training dataset.
The training dataset contains a larger number of interstitial defects, followed by substitutional and vacancy defects: this is mostly a consequence of there being a lot more options for intersitial and substitutional defects in terms of extrinsic species from across the periodic table, and also the longer time it takes for DFT optimization of these defect structures, leading to more intermediate geometries. This results in the comparatively higher RMSE for CFE prediction of vacancy defects (6.47 meV/atom) than substitutional (3.84 meV/atom) and interstitial (2.04 meV/atom) defects. The ALIGNN predictions for all types of structures are highly accurate with vanishingly small errors considering the total range of CFE values. GNN models for defect predictions reported in the recent literature 88, 43, 45 primarily focused on sampling defect configurations to train surrogate models, while we have adopted the approach of combining bulk and defect structures which enhances the overall generalizability and accuracy of the models.
Figure  3: Parity plots for ALIGNN models trained on the GGA dataset using (a) only bulk structures, (b) only defect structures, with predictions distinguished in terms of type of defect, and (c) both bulk and defect structures. (d) Active learning (AL) workflow implemented in this work: standard deviation of ALIGNN-predicted formation energy of novel defects is used to determine acquisition functions and identify the next set of DFT simulations to run. The ALIGNN model is then retrained with the new data and predictions are made for the remaining set of unexplored defects. (e) An ALIGNN vs DFT parity plot showing standard deviation in test set prediction across 100 separate models; these error bars are used to calculate the acquisition functions in the AL workflow.
Although state-of-the-art GNNs are very robust and powerful for modeling complex relationships within atomic structures 89, 87, 40, 90, 39, 91, 92, 93, their transferability beyond the trained dataset remains questionable. To evaluate this, we performed the following series of tests:
  1. 1.
  2. 2.
  3. 3.
ALIGNN shows poor transferability when trained only on bulk data and used to predict for defects; as shown in Figure S2(a), the prediction RMSEs range from $\sim$ 36 meV/atom for vacancies to >> 50 meV/atom for interstitial and substitutional defects. Since this model has not been exposed to specific configurations such as atomic relaxation around defect sites, it is unable to predict as accurately for defect structures as it does for bulk. Figure S2(a) also shows that ALIGNN uniformly under-predicts the CFE of all defect configurations, which could be attributed to the lower average CFE values in the bulk dataset compared to defects, as illustrated by the violin plot in Figure 2(e). The model trained only on interstitial defects does a reasonable job for vacancy defects, but the RMSE values are larger than from the models in Figure 3 and there are some very clear outliers, as pictured in Figure S2(b). Lastly, when the model is trained on only $\times2$$\times$2 supercell structures and used to predict for $\times3$$\times$3 supercells, the predictions show good accuracy but a slight tendency to over-estimate the CFE, hinting at the fact that ALIGNN may be capable of extrapolating across supercell sizes. These results suggest that the GNN models for CFE prediction could generalize across types of chemistries, structures, and system sizes for particular cases, but in general may need to be retrained and fine-tuned for specific datasets.
4 Expanded defect chemical space
The GNN models are ultimately intended for prediction and screening across thousands of possible novel defect configurations. Considering a series of Cd/Zn-Te/Se/S compositions, SQS structures, interfaces between them, and all possible native defects, group V dopants, impurities such as Cl and O, etc., there are potentially several hundred thousand single defects and double or triple defect complexes possible. Here, we defined an expanded defect chemical space using a subset of this entire available space, to keep the numbers tractable and enable predictions at a reasonable rate. We considered 9 possible compounds, namely CdTe, CdSe, ZnTe, three Cd-Se-Te compositions, and three Cd-Zn-Te compositions, and populated possible native defects (vacancies, self-interstitials, and anti-site substitutions), selected extrinsic defects (Cu, P, As, Sb, Cl, and O at Cd/Zn/Te/Se or interstitial sites), and selected defect complexes which are combinations of native defects and extrinsic defects. This chemical space is pictured in Figure S3 across the 9 compounds, showing a total of $\approx$ 25000 unique defect structures, accounting for both singular defects and defect complexes.
Symmetry-breaking is important for defect structures, and there could be many possible sites when considering defects in alloys or defect complexes. For all the defects described above, the Doped 86 package was used to identify possible defect sites in $\times3$$\times$3 supercells of binary and ternary compounds. For instance, AsTe (As at Te site) in Cd0.75Zn0.25Te has 108 non-equivalent Te sites, and all of them are taken into account in our expanded defect space. Theoretically, a non-symmetric supercell, such as any CdSexTe1-x compound, could host thousands of possible interstitial defects. Typically, interstitial defects are less stable than vacancies, substitutional defects, or anti-sites, as demonstrated in prior work  45. Therefore, we relaxed the criteria when generating interstitial defect configurations using Doped.
5 Active learning (AL) workflow with GNN-based geometry optimization
We employed an AL framework to efficiently explore the expanded defect space by iteratively training ensembles of 100 ALIGNN models each at different steps, before obtaining the best models pictured in Figure 3(a-c) and discussed earlier. The general AL workflow and results are pictured in Figure 3(d-e) and Figure S4. Mean and standard deviation of the CFE predictions were estimated using the ensemble of ALIGNN models. Based on a maximum uncertainty acquisition function, 200 of the most informative defect structures were selected for full DFT optimization and added to the training set in each AL cycle. Figure 3(e) shows an example model at one of the steps: plotted are ALIGNN-predicted CFE vs DFT-computed CFE (from subsequent calculations) for selected points from the expanded defect space, with error bars capturing the standard deviation in prediction across 100 models.
This iterative process rapidly improved model accuracy and predictive performance, even in relatively unexplored regions of the broader chemical space. Additional details of the AL workflow are presented in the SI. An important question we faced while making predictions for new defect structures is the following: how do we go beyond the CFE prediction for the initial defect configuration (including distorted structures generated from Doped) and determine the ground state configuration corresponding to the given defect? To achieve this, the ALIGNN prediction must be coupled with an optimization algorithm that relaxes atoms in the neighborhood of a defect and yields the lowest-energy structure. A version of this was presented in our past work 45, where hundreds of configurations with random atomic distortions were generated for any defect structure and GNN-predicted energies were used to determine the configuration closest to ground state.
  1. 1.
    Simulated Annealing (SA), which is a gradient-free stochastic method employing random atomic perturbations guided by an annealing schedule.
  2. 2.
    Bayesian Optimization (BO), which uses Gaussian processes to iteratively identify low energy configurations through atomic displacement exploration.
Figure  4: The DeFecT-FF workflow for simulating and understanding defects in Cd/Zn-Te/Se/S compounds. The training set includes bulk binary, ternary, and quaternary compounds, point-defect structures, grain boundary structures with defects, and interface structures with defects, simulated in five charge states ($q=+2$, +1+1, 0, 1-1, and 2-2) using both GGA-PBE and HSE06 functionals. Charge-specific MLFF models are trained to eventually enable geometry optimization of any new defect configurations.
Coupling ALIGNN prediction of CFE values with SA or BO enabled gradient-free energy minimization via atomic displacements applied within a cut-off radius around the defect center. These algorithms systematically searched for the lowest energy configuration using ALIGNN predictions at each step, allowing fast and guided optimization. As a case study, we applied both algorithms to optimize a Cd vacancy (VCd) defect in a $\times3$$\times$3 CdTe supercell, and the results are pictured in Figure  S5. Either method reaches a general energy convergence within 300 to 400 optimization steps which take a total of a few minutes to complete, but SA performs much better than BO: not only does it find a lower energy structure, but it actually discovers the configuration featuring a Te–Te dimer, which was reported by Kavanagh et al. 9 and which is easily missed by standard DFT optimization. To further evaluate the optimization capability of the ALIGNN model, we combined it with SA to optimize a variety of defects across the CdSexTe1-x and CdxZn1-xTe chemical space as shown in Figure  S6. For example, ClTe in Cd0.50Zn0.50Te was successfully optimized using ALIGNN+SA, as shown in Figure  S6 (a).
Figure  5: Performance of M3GNet-MLFF models trained on the GGA-PBE dataset, shown in terms of predicted vs DFT crystal formation energy parity plots. The models were trained separately for (a) defect configurations with charge q=+1, (b) neutral q=0 defect and bulk configurations, and (c) defect configurations with charge q=-1. Here, "bulk" refers to pristine supercells without any defects, "defect" means bulk supercells containing a point defect or defect complex, "interface" corresponds to defects located at CdTe-ZnTe interfaces, and "GB" indicates defects situated in CdTe grain boundary structures. Plots in (d–f) show the geometry optimization process taking into account 5 charge states for different example defect configurations: (d) a AsTe defect in CdTe, (e) a ZnCd defect at a CdTe/ZnTe interface, and (f) AsCd in a <111> CdTe grain boundary structure.
Figure  6: Statistics of the HSE06 dataset: number of bulk configurations corresponding to CdSexTe1-x, CdSxSe1-x, CdxZn1-xS, CdxZn1-xSe, CdxZn1-xTe, Cd0.5Zn0.5SxSe1-x, Cd0.5Zn0.5SexTe1-x, ZnSxSe1-x, and ZnSexTe1-x compositions; pie charts showing the number of defect configurations in different charge states in $\times2$$\times$2 supercells, $\times3$$\times$3 supercells, CdTe-ZnTe interfaces, and CdTe GB structures; and, example representations of point defects in interface and grain boundary structures.
Figure  7: Performance of the M3GNet-MLFF models trained on the augmented HSE06 dataset, shown in terms of predicted vs DFT crystal formation energy parity plots. The models were trained separately for (a) defect configurations with charge q=+1, (b) neutral q=0 defect and bulk configurations, and (c) defect configurations with charge q=-1. Here, "bulk" refers to pristine supercells without any defects, "defect" means bulk supercells containing a point defect or defect complex, "interface" corresponds to defects located at CdTe-ZnTe interfaces, and "GB" indicates defects situated at CdTe grain boundaries. Plots in (d–f) show the geometry optimization process taking into account 5 charge states for different example defect configurations: (d) an AsSe+ClSe defect complex in CdSe0.12Te0.88, (e) a Cdi defect at the CdTe/ZnTe interface, and (f) AsTe in <111> CdTe grain boundary structure.
Figure  8: The DeFecT-FF tool takes as input the bulk crystal, list of defects, bandgap, and VBM, performs MLFF-based geometry optimization, leverages an in-house chemical potential library to calculate charge-dependent energies $E(q)$ and defect formation energies $E_{f}$ as a function of the Fermi level $E_{F}$, and finally constructs $E_{f}$$E_{F}$ diagrams for defect thermodynamics. This tool is accessible via a nanoHUB web application.
6 Training crystal graph-based machine learning force field (MLFF)
M3GNet-based MLFF models were trained using the energies, atomic forces, and stresses extracted from GGA–PBE calculation trajectories. Figure 4 shows the workflow for training these models, eventually yielding the ability to optimize any new defect configuration and predict its ground state energy. Radial and three-body cutoffs were set to 6 Å and 6 Å, respectively. The loss was a weighted sum of RMSE for energies, forces, and stresses (weights = 1, 1, and 0.01, respectively). Training was performed on an NVIDIA A100 (80 GB) with batch size 64 and learning rate $5\times 10^{-4}$ until convergence. Geometry optimization with the MLFF used the FIRE algorithm in ASE 94, 95 with convergence criteria of mean atomic force $10^{-5}$ eV/Å or a maximum of 100 ionic steps. Models were trained separately for structures in five different charge states. To improve model accuracy on difficult configurations where predictions were poor, we used a two-stage training process:
Figure 5(a–c) show parity plots for the M3GNet-MLFF models trained for charge states $q=+1$, $q=0$, and q=1q=-1. Models for the $q=+2$ and q=2q=-2 charge states are presented in Figure S8. Each parity plot compares DFT-computed CFE for test set points with the corresponding values from the MLFF prediction for different types of structures: bulk (pristine supercells without defects), defects (bulk supercells containing a single point defect or defect complex), defects at CdTe-ZnTe interfaces, and defects in CdTe grain boundaries. Overall, the MLFF predictions show remarkably low RMSE values for all types of bulk and defect configurations, similar to the ALIGNN models.
For $q=0$, the test prediction RMSE for bulk structures is 8.6 meV/atom, a similarly low value of 6.8 meV/atom for defect structures, and 3.7 meV/atom for interface. Grain boundary defect structures show slightly larger errors of >> 9.1 meV/atom, which is expected given their more complex local environments and atomic rearrangements. Similar errors were seen for charged defect structures as well. The overall agreement between DFT and MLFF predictions remains generally strong, indicating the robustness of the model across diverse chemical and structural environments.
One of the major advantages of having an MLFF model rather than a direct energy prediction model is the ability to use predicted atomic forces to perform geometry optimization based on gradient-based energy minimization, which is more computationally efficient than gradient-free optimization as shown in Figure S9. For example, we optimized the AsTe+ClTe defect in CdSe0.12Te0.88 using both ALIGNN (direct energy) and M3GNet (MLFF) and found that M3GNet achieved the optimization at a substantially lower computational cost compared to ALIGNN. Figure 5(d–f) show three different examples of using the MLFF model for optimizing challenging defect configurations: an AsTe substitutional defect in CdTe, a ZnCd defect in a CdTe-ZnTe interface structure, and an AsCd substitutional defect in a CdTe GB structure. These cases represent chemically and structurally complex environments that are often found in devices with polycrystalline semiconductor thin films. The optimized q=+2,+1,0,1,2q=+2,+1,0,-1,-2 MLFF models were able to successfully capture local atomic rearrangements and produce low energy configurations consistent with DFT benchmarks. In each case, the MLFF-optimized configuration energy matches well with the DFT-optimized energy. Energy minimization is achieved in approximately 100 steps, with the entire relaxation process completing within a few minutes.
acronymacronymacronymacronymacronymacronymacronymacronymacronymacronymacronymacronymacronymacronym Dataset Supercell Size Data Points Bulk dataset from Cd/Zn-Te/Se/S binary compounds and alloys $\times2$$\times$2 5400 Defect dataset from 6 binary compounds $\times2$$\times$2 4302 (q=+2), 4201 (q=+1), 6203 (q=0), 4361 (q=-1), 4689 (q=-2) Defect dataset from CdSexTe1-x and CdxZn1-xTe alloys $\times3$$\times$3 371 (q=+2), 333 (q=+1), 402 (q=0), 321 (q=-1), 350 (q=-2) Defect dataset from CdTe/ZnTe interface $\times2$$\times$4 233 (q=+2), 223 (q=+1), 302 (q=0), 250 (q=-1), 270 (q=-2) Defect dataset from CdTe grain boundaries 133 (q=+2), 123 (q=+1), 163 (q=0), 121 (q=-1), 139 (q=-2)
Table 2: Number of data points (or structures) in the HSE06 dataset corresponding to different types of bulk or defect configurations, supercell sizes, and charge states.
@@ 7 MLFF models trained at hybrid functional accuracy
The ALIGNN and M3GNet-MLFF models trained on the GGA dataset can be enormously useful for optimizing new defect structures and screening for low energy defects. However, because of the well known limitations of the GGA functional in predicting electronic band edges and band gaps, chemical potentials, defect energetics, and defect levels 9, it is important to extend the DFT predictions to the hybrid HSE06 level of theory. Given the significant added computational expense of HSE06 calculations, it will take an extremely long time to generate HSE06 data for the upwards of 20,000 bulk and defect configurations in the GGA-PBE dataset. Thus, we performed HSE06 geometry optimization calculations on a subset of the structures on top of GGA optimization, which includes both initial and newly obtained configurations from the active learning schedule. Table 2 shows the number of structures of different types (bulk, defects, interfaces, GB) in different charge states and supercell sizes that eventually constituted the HSE06 dataset.
@@
HSE06 calculations for $3\times 3\times 3$ supercell defect structures with lattice constants around 20 Å were performed using $\Gamma$-point only 88, with a reduced plane-wave energy cutoff of 400 eV. The convergence thresholds for geometry optimization were set to $10^{-6}$ eV for energy and 0.01 eV Å-1 for forces. Calculations for grain boundary and interface structures were also performed using $\Gamma$-point only. Figure 6 shows the statistics of the compiled HSE06 dataset in terms of number of bulk Cd/Zn-Te/Se/S composition structures, and different types of defects in different charge states, including interface and GB defects. The HSE06 dataset represents 53.4% of the GGA dataset for bulk ($q=0$), and 63.8%, 71.6%, 79.5%, 72.4%, and 65.8% for defect charge states $q=+2$, $q=+1$, $q=0$, q=1q=-1, and q=2q=-2, respectively. Despite the smaller size, it remains well representative of the defect types and structural diversity of the entire chemical space. Violin plots showing the spread of CFE values in the HSE06 dataset are presented in Figure S10.

### v2

@@ 2 Description of the DFT Datasets
The entire Cd/Zn–Te/Se defect chemical space is extraordinarily large, as summarized in the Supporting Information (SI), Table SI. Treating every symmetry-inequivalent site in a $3\times 3\times 3$ (216-atom) cubic zincblende supercell results in thousands of possible defect configurations per composition, driven by the combinatorial explosion of native vacancies, self-interstitials, anti-site substitutions, eight types of extrinsic defects (dopants—Cu, As, P, N, Sb, Bi; impurities—Cl, O), and all unordered defect complexes, as illustrated in Figure S1. Even when restricted to only neutral charge states, the number of required DFT calculations grows to nearly $\approx$ 0.9 million for the entire chemical space. Thankfully, the defect structures collected from published literature and our prior work Mannodi-ZB, Mannodi_Toriyama_2020, Mannodi-Kanakkithodi2022-ck, Rahman_Gollapalli already represent a physically meaningful and chemically rich subset of this much larger defect chemical space. These literature-curated PBE-optimized structures span multiple compositions, charge states, and defect chemistries, providing a robust foundation for training the initial GNN models. We then employed an active-learning workflow to selectively launch new PBE calculations in regions of the chemical space where the ML model exhibited high uncertainty or sparse representation. Details of this dataset-building strategy are provided in the SI.
A carefully selected subset of PBE-optimized structures was further used to perform hybrid HSE06 calculations after adjusting the defect supercell lattice parameters to values from HSE06 optimization, before volume-fixed relaxation, as summarized in the Table SII. The resulting HSE06 dataset captures more than half of the structural and chemical diversity present in the full PBE dataset, and the statistical distribution of this HSE subset is illustrated in Figure 1. Importantly, all HSE calculations preserved every ionic-relaxation snapshot—not only the final minimum-energy configuration—yielding thousands of intermediate structures with their corresponding energies, forces, and stresses. This provides a significantly enriched dataset that captures full relaxation pathways rather than ground-state endpoints alone. For a complete statistical overview of the PBE and HSE datasets across all charge states, readers are referred to Table SIII, Table SIV and Table SV. Additional dataset visualization is provided in Figure S2, Figure S3, and Figure S4. In summary, our data generation pipeline proceeds as follows: literature data collection $\rightarrow$ initial PBE GNN model training $\rightarrow$ active-learning–driven expansion of the PBE dataset $\rightarrow$ HSE06 refinement of a subset of the PBE dataset.
Figure  2: (a–c) Parity plots comparing crystal formation energies from DFT and MLFF predictions for three representative charge states: (a) q = +1, (b) q = 0 (neutral), and (c) q = –1. The MLFF accurately reproduces the DFT energies with small errors, as indicated by the RMSE values shown in each panel. (d–f) Defect formation energies computed using MLFF predictions for a subset of the defect configurations shown in panels (a–c), compared against values from full DFT. The MLFF defect energies were obtained by adding DFT reference energies and applying charge corrections to the MLFF-predicted total energies.
Figure  3: Workflow for accelerated defect predictions using the DeFecT-FF framework. An initial defect structure (example: Asi in a mixed “2Te–2Se” local environment) is constructed and passed through the ShakeNBreak snb_1 symmetry-breaking procedure to generate a diverse set of competing defect geometries. These distorted configurations are rapidly relaxed using the rigorously optimized machine-learned force field to identify the lowest-energy structure prior to high-fidelity DFT calculations. The optimized geometry is then used to perform static HSE+SOC calculation, yielding accurate defect formation energy diagrams.
@@ 3 MLFF models trained at hybrid functional accuracy
We initially trained models on the PBE dataset using the ALIGNN framework ALIGNN and then employed an active learning strategy to launch new DFT calculations by targeting regions of the chemical space with largest prediction uncertainty. We then refined a representative subset of the PBE-optimized structures using the HSE06 functional and used these data to train an M3GNet-based m3gnet MLFF, using DFT-derived configurations, energies, forces, and stresses. Readers are referred to the SI for additional details on the ALIGNN training procedure and the active learning workflow used in this work. These discussions are supplemented by a series of figures in the SI (Figures S5–S13), which collectively illustrate the active learning pipeline, model transferability, ALIGNN-based structural optimization, comparisons with other MLFF models, and MLFF models on PBE dataset.
@@
HSE06 calculations were performed using $\Gamma$-point only Mosquera2, with a reduced plane-wave energy cutoff of 400 eV. The convergence thresholds for geometry optimization were set to $10^{-6}$ eV for energy and 0.01 eV Å-1 for forces. Figure 1 shows the statistics of the compiled HSE06 dataset in terms of the number of bulk Cd/Zn-Te/Se/S composition structures, and different types of defects. The HSE06 dataset represents 53.4% of the GGA dataset for bulk ($q=0$) structures, and 63.8%, 71.6%, 79.5%, 72.4%, and 65.8% for defect structures in charge states $q=+2$, $q=+1$, $q=0$, q=1q=-1, and q=2q=-2, respectively. Despite the smaller size, it remains well representative of the defect types and structural diversity of the entire chemical space. Violin plots showing the spread of the crystal formation energy (CFE, a per-atom energy difference between the crystal and its constituent atoms) values in the HSE06 dataset are presented in Figure S4.
Before training the MLFF models, we first evaluated the performance of state-of-the-art pretrained force-field frameworks, namely MACE mace, CHGNet chgnet and M3GNet m3gnet, by applying them directly to our PBE dataset. Although these models have demonstrated strong predictive capabilities on their training domains, they generalized poorly to the chemically diverse Cd/Zn–Te/Se/S systems investigated in this work as summarized in Table SVI. Root mean square error (RMSE) in CFE prediction ranged from 60 to 100 meV/atom, which it will turn out are much larger than errors from fine-tuned models. These shortcomings indicate that the pretrained models lack sufficient exposure to the chalcogenide defect chemical space considered here. Consequently, this motivated the development of a dataset-specific M3GNet-based MLFF trained on DFT-derived configurations, energies, forces, and stresses, enabling the level of accuracy required for reliable modeling of defect thermodynamics and structural relaxations.